A train travels 1 mile in 1 minute 30 seconds. At this rate, how many miles will it travel in two hours?
Answer by ptaylor(2198)

Distance(d)equals Rate(r) times Time(t) or d=rt; r=d/t and t=d/r
r=d/t
r=1/(1.5)=(2/3) mi/min
2hr =2*60=120 min
d=rt
d=(2/3)*120=80 mi
t=d/r=80/(2/3)=80*(3/2)=120 min
Hope this helps---ptaylor